# Flexible and efficient spatial extremes emulation via variational autoencoders

### v3

@@ Abstract
Emulating large spatial datasets is very valuable in producing a large ensemble of realistic replicates, quantifying uncertainty of various inference targets, and assessing risks associated with rare events such as joint high-threshold exceedances (Arpat and Caers, 2007). The efficacy of an emulator hinges greatly on its ability to characterize complex patterns of spatial variability in the studied phenomenon, especially the dependence structure in the tail. However, handling statistical computations for spatial models accommodating extremes remains a significant challenge for large datasets (Huser and Wadsworth, 2022).
Spatial emulation is particularly useful in climate science, where an abundance of simulations with high regional resolutions is often needed for detailed field-level insights and for uncertainty quantification (UQ). For example, watershed modeling (e.g., Sharma et al., 2021) requires a great number of simulations of heavy precipitation to assess regional riverine flood inundation hazards. In addition, numerous high-resolution sea surface temperature (SST) field replicates can aid in the identification of marine heatwaves (MHW) and the detection of coral reef zones prone to bleaching (Hughes et al., 2017). Current MHW detection methods, as described in Genevier et al. (2019), involve spatial averaging over predefined regions with extensive coral bleaching records, followed by calculating a daily percentile threshold using an 11-day window around the date of interest. To avoid relying on predefined subdomains or spatial averaging, we need a refined methodology to enhance the accuracy of spatial threshold exceedance estimations with better UQ.
Furthermore, emulators can be used as surrogate models to speed up and simplify mechanic-based deterministic models such as Earth system models (ESMs). These models typically rely on deterministic partial differential equations, featuring numerous parameters and requiring significant computational resources. In recent years, surrogate models have proven to produce realistic synthetic data with much less computational expense (see Gramacy, 2020, for an overview). They are also useful for inverse problems, model calibration and the specification of realistic model parameterizations for coupled processes.
@@
However, traditional methods for emulation such as those based on Gaussian processes (e.g., Gu et al., 2018), polynomial chaos expansions (e.g., Sargsyan, 2017), proper orthogonal decomposition (e.g., Iuliano and Quagliarella, 2013) and more recently, deep neural networks used in generative models such as generative adversarial networks (Goodfellow et al., 2014, GANs) and variational autoencoders (Kingma and Welling, 2013, VAEs), do not naturally accommodate extreme values, and certainly not dependent extreme values.
@@
Classical asymptotic max-stable processes exhibit dependence structures that are too rigid for environmental datasets and thus lead to overestimation of the spatial extent and intensity of extreme events (Huser et al., 2024). Recent spatial extremes models have addressed some of these limitations and offer more realistic tail dependence structures. Examples of such models include Gaussian scale mixtures (e.g., Opitz, 2016; Huser and Wadsworth, 2019) or $r$-Pareto processes (e.g., Thibaud and Opitz, 2015; de Fondeville and Davison, 2018), applied in the peaks-over-threshold framework, and max-infinitely divisible (max-id) models (e.g., Reich and Shaby, 2012; Padoan, 2013; Huser et al., 2021; Bopp et al., 2021; Zhong et al., 2022), applied in the block-maxima framework. However, these models often assume a stationary dependence structure across space and time, and the computational demands are significant even for moderately-sized datasets (up to about 500 locations; Zhang et al., 2022), limiting the applicability to high-resolution climate datasets.
In this work, we adopt a statistics-of-extremes perspective (Davison and Huser, 2015) and introduce a novel max-id model for spatial extremes with nonstationary dependence properties over space and time to explicitly accommodate concurrent and locally dependent extremes. We provide a detailed proof of these properties and propose embedding this flexible spatial extremes model within a variational autoencoder engine, referred to as the XVAE. This surrogate modeling framework facilitates rapid and efficient spatial modeling and data emulation in dimensions never explored before for dependent extremes. Moreover, we present a validation framework with wide applicability to assess emulation quality across low, moderate, and high values. A novel metric with theoretical guarantees is proposed, specifically tailored to evaluate the skill of a spatial model in fitting joint tail behavior.
The paper proceeds as follows. First, in Section 1, we detail how our novel max-id process can be embeded into the encoder-decoder VAE construction and study its highly flexible extremal dependence properties. Next, in Section 2, we introduce several model evaluation approaches to assess the emulation of spatial fields with dependent extremes. In Section 3, we validate the emulating power of our XVAE by simulation and demonstrate its advantages over a commonly used Gaussian process emulator. Then, in Section 4, we apply our XVAE to high-resolution Red Sea surface temperature extremes. Lastly, in Section 5, we conclude the paper and discuss perspectives for future work.
@@ 1 Methodology
This section provides background on VAEs, a presentation of our novel max-id model for spatial extremes, and a description of the construction of our XVAE. A detailed discussion on the tail properties of the model is also included. Hereafter, we shall denote random variables with capital letters and fixed or observed quantities with lowercase letters.
1.1 Variational autoencoder background
Bayesian hierarchical models characterized by a lower-dimensional latent process are amenable to use of the variational autoencoder for inference and generative simulation. Specifically, consider the joint distribution
where $\boldsymbol{x}$ and $\boldsymbol{z}$ are observations of a (e.g., physical) process $\boldsymbol{X}\in\mathbb{R}^{n_{s}}$ and realizations of a latent process $\boldsymbol{Z}\in\mathbb{R}^{K}$, respectively, and $K\ll n_{s}$. The data model $p_{\boldsymbol{\theta}}(\boldsymbol{x}\mid\boldsymbol{z})$ and the latent process model $p_{\boldsymbol{\theta}}(\boldsymbol{z})$ may depend on different subsets of parameters $\boldsymbol{\theta}$. In the case of spatial data, the vector $\boldsymbol{X}$ may be observations of a spatial process $\{X(\boldsymbol{s}):\boldsymbol{s}\in\mathcal{S}\}$ at $n_{s}$ locations, and $\boldsymbol{Z}$ may be random coefficients from a low-rank basis expansion representation of $\boldsymbol{X}$. To emulate an observed high-dimensional data $\boldsymbol{x}$ (i.e., generate new realizations of $\boldsymbol{X}$), an ideal probabilistic framework would be to: (1) obtain a good estimate $\hat{\boldsymbol{\theta}}$ of the parameters $\boldsymbol{\theta}$ given the observations $\boldsymbol{x}$ and generate independent (latent) variables $\boldsymbol{Z}^{1},\ldots,\boldsymbol{Z}^{L}$ from the posterior density $p_{\hat{\boldsymbol{\theta}}}(\boldsymbol{z}\mid\boldsymbol{x})$; (2) generate $\boldsymbol{X}^{l}$ from the data model (posterior predictive distributions) $p_{\hat{\boldsymbol{\theta}}}(\boldsymbol{x}\mid\boldsymbol{Z}^{l})$, $l=1,\ldots,L$. If the characterization of the distributions is reasonable, the resulting emulated replicates $\{\boldsymbol{X}^{1},\ldots,\boldsymbol{X}^{L}\}$ should resemble the original input $\boldsymbol{x}$, with meaningful variations among the emulated samples.
Unfortunately, the posterior density $p_{\boldsymbol{\theta}}(\boldsymbol{z}\mid\boldsymbol{x})$ is often intractable because the marginal likelihood $p_{\boldsymbol{\theta}}(\boldsymbol{x})=\int p_{\boldsymbol{\theta}}(\boldsymbol{x},\boldsymbol{z}){\rm d}\boldsymbol{z}$, a $K$-dimensional integral, typically does not have an analytical form. This makes the maximum likelihood estimation of the parameters $\boldsymbol{\theta}$ difficult even for a moderately complicated data model $p_{\boldsymbol{\theta}}(\boldsymbol{x}\mid\boldsymbol{z})$. Imposing another prior model on $\boldsymbol{\theta}$ and implementing a Markov chain Monte Carlo (MCMC) algorithm to draw samples of $(\boldsymbol{Z},\boldsymbol{\theta})$ simultaneously can be a possible solution; the draws of $\boldsymbol{Z}$ from the Markov chain are treated as the samples from $p_{\boldsymbol{\theta}}(\boldsymbol{z}\mid\boldsymbol{x})$ which can then be fed in $p_{\boldsymbol{\theta}}(\boldsymbol{x}\mid\boldsymbol{z})$ to generate new $\boldsymbol{X}$. But MCMC can be computationally intensive when the likelihood $p_{\boldsymbol{\theta}}(\boldsymbol{x},\boldsymbol{z})$ is expensive to evaluate, and algorithm tuning can be challenging when the dimension of $(\boldsymbol{Z},\boldsymbol{\theta})$ is large, particularly in the case of complex spatial and/or temporal dependence in the data.
In variational Bayesian inference, the true posterior $p_{\boldsymbol{\theta}}(\boldsymbol{z}\mid\boldsymbol{x})$ is approximated by a so-called variational distribution that is relatively easy to evaluate. The variational autoencoder (VAE) proposed by Kingma and Welling (2013) formulates the variational distribution (which they call a probabilistic encoder) as a multilayer perceptron (MLP) neural network denoted by $q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}\mid\boldsymbol{x})$, in which $\boldsymbol{x}$ is the input, $\boldsymbol{z}$ is the output and $\boldsymbol{\phi}_{e}$ are the so-called weights and biases of the neural network. This construction greatly facilitates fast non-linear encoding of a high-dimensional process $\boldsymbol{X}$ to lower-dimensional $\boldsymbol{Z}$ in the latent space. Next, to generate emulated data replicates from $p_{\boldsymbol{\theta}}(\boldsymbol{x}\mid\boldsymbol{Z})$, Kingma and Welling (2013) use another MLP neural network (which they call a probabilistic decoder) acting as an estimator for the model parameters $\boldsymbol{\theta}$. Denoting the weights and biases of the decoder by $\boldsymbol{\phi}_{d}$, we write $\hat{\boldsymbol{\theta}}_{\text{NN}}=\mathrm{DecoderNeuralNet}_{\boldsymbol{\phi}_{d}}(\boldsymbol{z})$ and denote through an abuse of notation
$p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x},\boldsymbol{z})\equiv p_{\hat{\boldsymbol{\theta}}_{\text{NN}}}(\boldsymbol{x},\boldsymbol{z}),\quad p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x})\equiv\int p_{\hat{\boldsymbol{\theta}}_{\text{NN}}}(\boldsymbol{x},\boldsymbol{z})\text{d}\boldsymbol{z},$
and $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{z}\mid\boldsymbol{x})=p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x},\boldsymbol{z})/p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x})$.
The encoding parameters $\boldsymbol{\phi}_{e}$ and decoding parameters $\boldsymbol{\phi}_{d}$ can be estimated by maximizing the evidence lower bound (ELBO), which is defined by
$\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}(\boldsymbol{x})=\log p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x})-D_{KL}\left\{q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}\mid\boldsymbol{x})\;||\;p_{\boldsymbol{\phi}_{d}}(\boldsymbol{z}\mid\boldsymbol{x})\right\},$ (1)
in which $\log p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x})$ is called the evidence for $\boldsymbol{x}$ under $\hat{\boldsymbol{\theta}}_{\text{NN}}$, and the second term in (1) is the Kullback–Leibler (KL) divergence between $q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}\mid\boldsymbol{x})$ and $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{z}\mid\boldsymbol{x})$, which is non-negative. Therefore, maximizing the ELBO with respect to $\boldsymbol{\phi}_{e}$ and $\boldsymbol{\phi}_{d}$ is equivalent to maximizing the log-likelihood $\log p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x})$ while minimizing the difference between the approximate posterior density $q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}\mid\boldsymbol{x})$ and the true posterior density $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{z}\mid\boldsymbol{x})$ (under $\hat{\boldsymbol{\theta}}_{\text{NN}}$).
Therefore, the ELBO can be approximated by Monte Carlo as
$\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}(\boldsymbol{x})\approx\frac{1}{L}\sum_{l=1}^{L}\log\frac{p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x},\boldsymbol{Z}^{l})}{q_{\boldsymbol{\phi}_{e}}(\boldsymbol{Z}^{l}\mid\boldsymbol{x})},$ (2)
where $\boldsymbol{Z}^{1},\ldots,\boldsymbol{Z}^{L}$ are independent draws from $q_{\boldsymbol{\phi}_{e}}(\cdot\mid\boldsymbol{x})$. If there are replicates of the process, $\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{n_{t}}$, then $\sum_{t=1}^{n_{t}}\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}(\boldsymbol{x}_{t})$ is considered. In traditional VAEs (e.g., Kingma et al., 2019; Cartwright et al., 2023), Gaussianity is assumed for both the data model $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x}_{t}\mid\boldsymbol{z}_{t})$ and the encoder $q_{\phi_{e}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$, with the prior $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{z}_{t})$ often set as a simple multivariate normal distribution $N(\boldsymbol{0},\boldsymbol{I}_{K})$. However, the Gaussian assumptions limit the VAE’s ability to capture heavy-tailed distributions and intricate dependence structures.
1.2 A nonstationary max-id model for VAE-based emulation
@@
To better emulate spatial data with extremes, we define $p_{\boldsymbol{\theta}}(\boldsymbol{x}\mid\boldsymbol{z})$ indirectly through the construction of a novel flexible nonstationary spatial extremes model that can easily be incorporated into a VAE. The flexible spatial extremes model is described in Section 1.2.1, followed by a detailed description the XVAE in Section 1.2.2. Note that implementation details related to the XVAE neural networks are provided in the Supplementary Material.
1.2.1 Flexible nonstationary max-id spatial extremes model

### v4

# Flexible and efficient emulation of spatial extremes processes via variational autoencoders

@@ Abstract
1 Introduction
Statistical emulators, pioneered by Sacks et al. (1989) and Kennedy and O’Hagan (2001), have been mostly used to accurately approximate patterns and relationships in deterministic model outputs (e.g., from climate models, fluid dynamics or other physical systems), which are computationally prohibitive to obtain at high spatio-temporal resolution. Statistical emulators, used as surrogate models, have thus been beneficial for model calibration, where one estimates unknown parameters of a deterministic model by aligning model outputs with observed data (e.g., Higdon et al., 2004; Bayarri et al., 2007; Chang et al., 2016; Gopalan and Wikle, 2022).
Another related key application of statistical emulators is to use them with real or model-output data to quickly generate large ensembles of realistic simulations of complex (random or deterministic) spatio-temporal processes. This is especially advantageous to improve uncertainty quantification (UQ) for various inference targets (see, e.g., Gramacy, 2020), particularly when assessing risks related to rare events, e.g., defined as joint exceedances over high thresholds. For instance, current marine heatwave (MHW) detection methods often involve calculating percentile thresholds empirically from a quite limited number of daily sea surface temperature (SST) observations, averaged spatially over relatively coarse regions (Hughes et al., 2017). Emulating SST data over space and time can thus enhance the estimation of, and UQ for, extreme hotspots defined as regions experiencing high temperatures simultaneously. We come back to such an application in Section 5.
@@ 1 Introduction
The efficacy of an emulator hinges greatly on its ability to capture complex spatial variability, which is particularly true when interest lies in the tail dependence structure. However, traditional emulation methods—such as those based on Gaussian processes (e.g., Gu et al., 2018), polynomial chaos expansions (e.g., Sargsyan, 2017) and more recently, deep neural networks such as generative adversarial networks (Goodfellow et al., 2014, GANs) and variational autoencoders (Kingma and Welling, 2013, VAEs)—do not naturally accommodate nor realistically reproduce extreme values, and certainly not dependent extremes. By contrast, classical spatial models justified by extreme-value theory are often overly computationally costly to fit with large datasets (Huser and Wadsworth, 2022).
The main methodological contributions of this work are threefold. First, we introduce a novel max-infinitely divisible (max-id) model for spatial extremes with nonstationary dependence structure that varies over both space and time, and formally prove that it flexibly accommodates concurrent and locally dependent extremes. Second, we propose embedding this complex spatial extremes model within a VAE engine, referred to as the XVAE, to facilitate fast inference and simulation in high dimensions. Third, we develop a general validation framework to assess an emulator’s quality across low, moderate, and high values. A novel metric with theoretical guarantees is proposed, specifically tailored to evaluate the skill of a spatial model in reproducing dependent extremes. Given that most validation approaches either lack emphasis on the joint tail behavior (e.g., Gneiting and Raftery, 2007) or rely on simple bivariate summaries, our proposed framework is a valuable additional tool that complements standard model validation techniques.
The paper is organized as follows: In Section 2, we concisely review classical spatial extremes models and VAEs. In Section 3, we detail our novel max-id process, derive its flexible extremal dependence properties, and demonstrate how to integrate it within a VAE. We also present our general model validation framework to evaluate spatial process emulators with an emphasis on dependent extremes. In Section 4, we validate the emulating power of our XVAE through simulations, and compare it to a Gaussian process emulator. In Section 5, we apply the XVAE to high-resolution Red Sea SST, and use it to efficiently enhance UQ of extreme sea temperature hotspot estimates. Finally, in Section 6, we conclude with some discussion on future research directions.
2 Background
This section provides background on spatial extremes models and VAEs. Random variables are denoted with capital letters and fixed or observed quantities with lowercase letters.
2.1 Spatial extremes modeling
In the spatial extremes literature, extremal dependence is commonly measured by
$\chi_{ij}(u)=\Pr\{F_{j}(X_{j})>u\mid F_{i}(X_{i})>u\}=\frac{\Pr\{F_{j}(X_{j})>u,F_{i}(X_{i})>u\}}{1-u}\in[0,1],$ (1)
for some threshold $u\in(0,1)$ and where $F_{i}$ and $F_{j}$ are continuous marginal distribution functions for the random variables $X_{i}$ and $X_{j}$, respectively. When $u\approx 1$, $\chi_{ij}(u)$ quantifies the probability that one variable is extreme given that another variable is similarly extreme. If $\chi_{ij}=\lim_{u\rightarrow 1}\chi_{ij}(u)=0$, $X_{i}$ and $X_{j}$ are said to be asymptotically independent (AI), and if $\chi_{ij}=\lim_{u\rightarrow 1}\chi_{ij}(u)>0$, $X_{i}$ and $X_{j}$ are asymptotically dependent (AD).
Classical asymptotic models such as max-stable (Davison and Huser, 2015; Davison et al., 2012, 2019) or generalized Pareto (Ferreira and de Haan, 2014; Thibaud and Opitz, 2015; de Fondeville and Davison, 2018) processes always have $\chi_{ij}>0$, unless $X_{i}$ and $X_{j}$ are exactly independent. Conversely, Gaussian processes—or marginal transformations thereof—always have $\chi_{ij}=0$, unless $X_{i}$ and $X_{j}$ are perfectly dependent.
In practice, extremal dependence (i.e., $\chi_{ij}(u)$) estimated from environmental processes is often observed to decay as events get more extreme (i.e., $u\to 1$) and to become spatially more localized as their intensity increases (Huser et al., 2024). This phenomenon was observed in numerous studies, e.g., on Dutch wind gust maxima (Huser et al., 2021), threshold exceedances of the daily Fosberg fire index (Zhang et al., 2022), and winter maximum precipitation data over the Midwest of the U.S. (Zhang et al., 2023), just to name a few examples. This implies that the stability property of max-stable and generalized Pareto models is often physically inappropriate. However, a weakening $\chi_{ij}(u)$ as $u$ increases does not necessarily lead to AI, and Gaussian processes have a quite restrictive tail behavior. Therefore, we seek to develop models that exhibit much more flexible tail characteristics and that do not assume an extremal dependence class a priori. This is especially important for risk assessment when extrapolating beyond the observed data, as misspecifying the extremal dependence regime can lead to grossly inaccurate joint tail probability estimates.
@@ 2.1 Spatial extremes modeling
Recent spatial extremes models have addressed some of these limitations and offer more realistic tail properties. Examples of such models include random scale mixtures (e.g., Opitz, 2016; Huser et al., 2017; Huser and Wadsworth, 2019), usually applied in the peaks-over-threshold framework, and max-id models (e.g., Reich and Shaby, 2012; Padoan, 2013; Huser et al., 2021; Bopp et al., 2021; Zhong et al., 2022), mostly applied in the block-maxima framework; see Huser and Wadsworth (2022) for an overview. However, these models often assume a stationary dependence structure (in particular, the same dependence class at fixed space-time lag) across space and time, and do not always represent long-range dependence realistically over large geographical domains (Hazra et al., 2024). Moreover, the computational demands for fitting such models using standard inference techniques are significant even for moderately-sized datasets (see, e.g., Zhang et al., 2022, who apply such a model on 93 sites), hampering their applicability to high-resolution climate datasets.
More recently, several attempts have been made to exploit advances in deep learning to facilitate the modeling, inference, and simulation of multivariate and spatial extremes. Richards and Huser (2022) and Pasche and Engelke (2024) use deep extreme quantile regression models to improve the modeling of marginal extremes in spatial and temporal settings, respectively. Boulaguiem et al. (2022) use a deep convolutional GAN (called extGAN) to learn the dependence structure of spatial extremes; their approach, however, does not impose any parametric constraint on the extremal dependence structure, which leads to AI. By contrast, Lafon et al. (2023) develop a VAE tailored to multivariate regularly varying (i.e., jointly heavy-tailed) data; their approach thus only applies to AD data, and it has so far only been validated in small dimensions (specifically, 5 sites in their application). Lenzi et al. (2023), Sainsbury-Dale et al. (2024, 2023) and Richards et al. (2023) use deep learning methods for fast likelihood-free inference with parametric spatial extreme-value models. These inference methods are amortized (Zammit-Mangion et al., 2025) in the sense that they are very fast after an initial upfront computational cost has been incurred to train a neural network. Such methods are simulation-based and cannot, however, easily handle highly-parameterized models such as nonstationary processes (but see Zammit-Mangion and Wikle, 2020); further, they are meant to provide parameter inferences, not to simultaneously generate realistic data simulations. In the same vein, Majumder et al. (2024) use deep learning to speed up updates in a Markov chain Monte Carlo (MCMC) algorithm, in order to fit a complex, but stationary, spatial dependence model.
In this work, we aim to develop the first VAE able to emulate high-resolution, non-stationary, spatio-temporal extremes data, and that can provide fast parameter inferences and UQ, along with realistic data simulations accounting for the possibility of AI and AD.
2.2 Variational autoencoder background
Bayesian hierarchical models with a lower-dimensional latent process can leverage VAEs for inference and statistical emulation. These models typically assume the joint distribution
where $\boldsymbol{x}$ represents observations of a (e.g., physical) process $\boldsymbol{X}\in\mathbb{R}^{n_{s}}$ and $\boldsymbol{z}$ denotes realizations of a latent process $\boldsymbol{Z}\in\mathbb{R}^{K}$. In the case of spatial data, the vector $\boldsymbol{X}$ may be observations of a spatial process $\{X(\boldsymbol{s}):\boldsymbol{s}\in\mathcal{S}\}$ at $n_{s}$ locations, and $\boldsymbol{Z}$ may be random coefficients from a low-rank basis expansion representation of $\boldsymbol{X}$.
An ideal probabilistic framework for emulating an observed $\boldsymbol{x}$ (a number of times, $L$, say) would be to: (1) estimate parameters $\hat{\boldsymbol{\theta}}$ given the input $\boldsymbol{x}$ and sample latent variables $\boldsymbol{Z}^{1},\ldots,\boldsymbol{Z}^{L}$ from the posterior $p_{\hat{\boldsymbol{\theta}}}(\boldsymbol{z}\mid\boldsymbol{x})$; (2) generate $\boldsymbol{X}^{l}$ from the posterior predictive distributions $p_{\hat{\boldsymbol{\theta}}}(\boldsymbol{x}\mid\boldsymbol{Z}^{l})$, $l=1,\ldots,L$. If the characterization of the distributions is reasonable, the new realizations $\{\boldsymbol{X}^{1},\ldots,\boldsymbol{X}^{L}\}$ should resemble the input $\boldsymbol{x}$, with meaningful variations among the replicates. However, the posterior $p_{\boldsymbol{\theta}}(\boldsymbol{z}\mid\boldsymbol{x})$ is often intractable when the marginal likelihood $p_{\boldsymbol{\theta}}(\boldsymbol{x})=\int p_{\boldsymbol{\theta}}(\boldsymbol{x},\boldsymbol{z}){\rm d}\boldsymbol{z}$ does not have an analytical form, complicating parameter estimation for high-dimensional data using methods like MCMC.
Under the variational Bayes framework, the VAE proposed by Kingma and Welling (2013) approximates the posterior $p_{\boldsymbol{\theta}}(\boldsymbol{z}\mid\boldsymbol{x})$ using a so-called probabilistic encoder. Formulated via a multilayer perceptron (MLP) neural network, the encoder maps the input $\boldsymbol{x}$ to a variational distribution in the latent space denoted by $q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}\mid\boldsymbol{x})$, in which $\boldsymbol{\phi}_{e}$ are the weights and biases of the encoder network. Then, a sample $\boldsymbol{Z}$ from ${q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}\mid\boldsymbol{x})}$ is generated and a decoder network acts as an estimator for the model parameters: $\hat{\boldsymbol{\theta}}_{\text{NN}}=\mathrm{DecoderNeuralNet}_{\boldsymbol{\phi}_{d}}(\boldsymbol{Z})$. Finally new realizations of $\boldsymbol{X}$ can be generated from $p_{\hat{\boldsymbol{\theta}}_{\text{NN}}}(\boldsymbol{x}\mid\boldsymbol{Z})$.
We denote through an abuse of notation $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x},\boldsymbol{z})\equiv p_{\hat{\boldsymbol{\theta}}_{\text{NN}}}(\boldsymbol{x}\mid\boldsymbol{z})p_{\hat{\boldsymbol{\theta}}_{\text{NN}}}(\boldsymbol{z})$, $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x})=\int p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x},\boldsymbol{z})\text{d}\boldsymbol{z}$ and $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{z}\mid\boldsymbol{x})=p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x},\boldsymbol{z})/p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x})$. The VAE is typically trained by maximizing the evidence lower bound (ELBO), which balances the log-likelihood and the Kullback–Leibler (KL) divergence between the approximate and true posteriors:
$\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}(\boldsymbol{x})=\log p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x})-D_{KL}\left\{q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}\mid\boldsymbol{x})\;||\;p_{\boldsymbol{\phi}_{d}}(\boldsymbol{z}\mid\boldsymbol{x})\right\}.$ (2)
Here, $\log p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x})$ is called the evidence for $\boldsymbol{x}$, and the KL divergence is non-negative.
In traditional VAEs (e.g., Kingma et al., 2019; Cartwright et al., 2023), Gaussianity is assumed for both the data model $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x}\mid\boldsymbol{z})$ and the encoder $q_{\phi_{e}}(\boldsymbol{z}\mid\boldsymbol{x})$, with the prior $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{z})$ often set as a multivariate normal distribution. However, such Gaussian assumptions limit the VAE’s ability to capture heavy-tailed distributions (Lafon et al., 2023).
@@ 3 Methodology
To better emulate spatial data with extremes, we define $p_{\boldsymbol{\theta}}(\boldsymbol{x}\mid\boldsymbol{z})$ indirectly through the construction of a novel flexible nonstationary spatial extremes model, introduced in Section 3.1. A detailed description on how the model is integrated into the XVAE is given in Section 3.2. In Section 3.3, we propose a new validation framework that is tailored to assess skill in fitting both the full range and the joint tail behavior of model outputs.
3.1 Flexible nonstationary max-id spatial extremes model